# Hybrid Meta-learners for Estimating Heterogeneous Treatment Effects

### v1

@@ Appendix C Additional Results on ACIC 2016
T-learner 1.99±.05plus-or-minus1.99.051.$9$05 $2.06\pm.06$ $2.16\pm.05$ $2.30\pm.06$
TARNet $1.56\pm.05$ $1.65\pm.05$ 1.77±.05plus-or-minus1.77.051.$7$05 $1.90\pm.05$
@@
DR-learner $1.61\pm.05$ $1.72\pm.06$ 1.66±.05plus-or-minus1.66.051.$6$05 1.77±.06plus-or-minus1.77.061.$7$06
@@
TARNet-WR 1.88±.06plus-or-minus1.88.061.$8$06 $1.96\pm.06$ $1.91\pm.06$ 1.99±.06plus-or-minus1.99.061.$9$06
OffsetNet $2.05\pm.07$ 2.11±.07plus-or-minus2.11.072.$1$07 $2.15\pm.07$ $2.20\pm.08$
FlexTENet $1.79\pm.06$ 1.88±.07plus-or-minus1.88.071.$8$07 $1.83\pm.06$ $1.92\pm.06$
@@
H-learner (X) $\bm{1.55\pm.05}$ 1.66±.06plus-or-minus1.66.061.$6$06 $\bm{1.58\pm.05}$ $\bm{1.68\pm.05}$
@@
T-learner $2.09\pm.05$ $2.21\pm.06$ $2.07\pm.05$ $2.17\pm.05$
TARNet 1.66±.05plus-or-minus1.66.051.$6$05 $1.78\pm.06$ 1.66±.05plus-or-minus1.66.051.$6$05 $1.76\pm.05$

### v2

@@ Appendix D Additional Results on ACIC 2016
Table 5: Performance on the ACIC 2016 benchmark dataset across different heterogeneity levels.
@@
T-learner $1.99\pm.05$ $2.06\pm.06$ $2.16\pm.05$ $2.30\pm.06$
TARNet $1.56\pm.05$ $1.65\pm.05$ $1.77\pm.05$ $1.90\pm.05$
@@
Table 6: Performance on the ACIC 2016 benchmark dataset across different confounding levels.
@@
DR-learner $1.61\pm.05$ $1.72\pm.06$ $1.66\pm.05$ $1.77\pm.06$
@@
TARNet-WR $1.88\pm.06$ $1.96\pm.06$ $1.91\pm.06$ $1.99\pm.06$
OffsetNet $2.05\pm.07$ $2.11\pm.07$ $15\pm.07 2.2$$20\pm.08$
FlexTENet $1.79\pm.06$ $1.88\pm.07$ $1.83\pm.06$ $1.92\pm.06$
@@
H-learner (X) $\bm{1.55\pm.05}$ $1.66\pm.06$ $\bm{1.58\pm.05}$ $\bm{1.68\pm.05}$
Table 7: Performance on the ACIC 2016 benchmark dataset across different overlap levels.
@@
T-learner $09\pm.05 2.2$$21\pm.06$ $2.07\pm.05$ $2.17\pm.05$
TARNet $1.66\pm.05$ $1.78\pm.06$ $1.66\pm.05$ $1.76\pm.05$